OpenStudy (jalsto6761):

five added to twice Erik's age is the same as 3 times his age minus 2 how old is Erik?

OpenStudy (gunsodo):

Let his age be x. Could you write the equation?

OpenStudy (jalsto6761):

no that is what i need help on

OpenStudy (mathmale):

If Eric's age is x, what is "twice Eric's age?"

OpenStudy (miasan5653):

okay. 5 + 2x because you add five to twice his age. Based on that do you think you can do the second part

OpenStudy (jalsto6761):

so its 5+ 2x =3

OpenStudy (mathmale):

Not quite. What is Eric's age minus 2?

OpenStudy (miasan5653):

three times his age minus two. not just three

OpenStudy (jalsto6761):

5+2x-2=3

OpenStudy (miasan5653):

5+2x will equal 3x-2 because these are each part of the two separate equations

OpenStudy (miasan5653):

To get the actual age, set them equal to each other and solve for x

OpenStudy (jalsto6761):

so the equation is 5+2x=3x-2

OpenStudy (miasan5653):

Yep

OpenStudy (miasan5653):

You are going to get the 5 and --2 on one side by adding 2 to 5 the other side will have 3x-2x

OpenStudy (jalsto6761):

so would you do 2x+3x then do -2+5

OpenStudy (miasan5653):

exactly

OpenStudy (miasan5653):

Wait, I shouldn't have said exactly. I'll do a drawing to make it more clear,

OpenStudy (miasan5653):



OpenStudy (jalsto6761):

don't you box in the virable and turn 5 into a -5

OpenStudy (miasan5653):

Nope. The main goal here is to isolate x and you can do that by moving all of the xs to the right. The five will not move at all and only gets added to.

OpenStudy (miasan5653):

Close the question when you are done @jalsto6761

OpenStudy (mathmale):

If you believe x=7 may be correct, then please check your work thru substitution. The problem states: "five added to twice Erik's age is the same as 3 times his age minus 2 how old is Erik?" What is twice erik's age? What is five added to twice Eric's age? What is 3 times Eric's age? What is 3 times his age, minus 2? Make the appropriate substitutions in the pertinent equation. is the equation now true or false?

OpenStudy (miasan5653):

This is correct right? I didn't mess up somewhere? @mathmale